OpenStudy (anonymous):

Graph the circle (x + 2)2 + (y + 1)2 = 32. Label the center and at least four points on the circle.

OpenStudy (tkhunny):

Can you find the Center?

OpenStudy (anonymous):

(-2,-1), i just dont get how to find the 4 points b/c 32 squared is not a whole number

OpenStudy (zale101):

the square root of 32 \[32=\sqrt{32}= \sqrt{16*2}=4\sqrt{2}\]

OpenStudy (anonymous):

but how do i plot that on a graph? that what confuessing me

OpenStudy (anonymous):

@Zale101?

OpenStudy (zale101):

calculate the expression \[4\sqrt{2}\]

OpenStudy (anonymous):

5.657

OpenStudy (zale101):

\[4\sqrt{2}\] is most definitely a number

OpenStudy (zale101):

yes

OpenStudy (anonymous):

so do i keep it 5.6 or should i round it to 6? for graphing reason

OpenStudy (zale101):

round the 6

OpenStudy (anonymous):

ok thank you!!!

OpenStudy (zale101):

np